# Deep Reinforcement Learning based Autonomous Decision-Making for Cooperative UAVs: A Search and Rescue Real World Application

### v1

@@ Abstract
This paper proposes a holistic framework for autonomous guidance, navigation, and task distribution among multi-drone systems operating in Global Navigation Satellite System (GNSS)-denied indoor settings. We advocate for a Deep Reinforcement Learning (DRL)-based guidance mechanism, utilising the Twin Delayed Deep Deterministic Policy Gradient algorithm. To improve the efficiency of the training process, we incorporate an Artificial Potential Field (APF)-based reward structure, enabling the agent to refine its movements, thereby promoting smoother paths and enhanced obstacle avoidance in indoor contexts. Furthermore, we tackle the issue of task distribution among cooperative UAVs through a DRL-trained Graph Convolutional Network (GCN). This GCN adeptly represents the interactions between drones and tasks, facilitating dynamic and real-time task allocation that reflects the current environmental conditions and the capabilities of the drones. Such an approach fosters effective coordination and collaboration among multiple drones during search and rescue operations or other exploratory endeavours. Lastly, to ensure precise odometry in environments lacking GNSS, we employ Light Detection And Ranging Simultaneous Localisation and Mapping complemented by a depth camera to mitigate the hallway problem. This integration offers robust localisation and mapping functionalities, thereby enhancing the system’s dependability in indoor navigation. The proposed multi-drone framework not only elevates individual navigation capabilities but also optimises coordinated task allocation in complex, obstacle-laden environments. Experimental evaluations conducted in a setup tailored to meet the requirements of the NATO Sapience Autonomous Cooperative Drone Competition demonstrate the efficacy of the proposed system, yielding outstanding results and culminating in a first-place finish in the 2024 Sapience competition.
@@ II-A Deep Reinforcement Learning for Autonomous Guidance and Obstacle Avoidance
To address this inefficiency, researchers have incorporated expert knowledge into DRL training, enhancing performance through pre-training. Techniques such as Deep Q Learning from Demonstrations (DQfD) [12] and Deep Deterministic Policy Gradients from Demonstrations (DDPGfD) [50] leverage expert-labeled data to pre-train networks, reducing reliance on extensive reward engineering while preserving the benefits of DRL.
Non-Expert Policy Enhanced DRL (NPE-DRL) [54] and DDPG-APF [13] leverage the Artificial Potential Field (APF) to generate non-expert actions that guide the agent during training. Both methods gradually reduce the influence of APF to allow for exploration and optimal policy discovery. NPE-DRL operates in a discrete action space, while DDPG-APF utilises a continuous action space. Each approach demonstrates improved performance and sample efficiency, enabling the agent to converge on effective solutions with fewer training samples.
This paper builds by integrating the APF directly into the reward structure, rather than using it as a separate guiding mechanism. This approach is expected to provide tighter control over the UAV’s policy, especially within the more complex indoor environments considered here, which extends beyond the scenarios addressed by NPE-DRL and DDPG-APF.
@@ III-A1 Twin Delayed Deep Deterministic Policy Gradient
TD3 [8] builds on the DDPG [23] algorithm to improve stability in continuous control tasks, addressing common issues like overestimation bias. TD3 follows an actor-critic structure where the actor network outputs actions and the critic networks evaluate the expected returns. TD3 introduces three core innovations:
  1. 1.
    Clipped Double Q-Learning: Two Q-networks are used, with updates based on the minimum Q-value, reducing overestimation.
  2. 2.
    Delayed Policy Updates: The actor is updated less frequently than the critics, allowing the critics to stabilise and prevent premature policy updates.
  3. 3.
    Target Policy Smoothing: This is a regularisation strategy where the target actions have a small noise signal (smoothing) added before being passed to the target critic networks. This variation prevents overfitting of the Q-networks to small peaks in the value estimate. This smoothing of the target actions leads to a lower variation in the Q targets and lowers the chance of functional approximation error.
The training process involves the agent observing the environment, selecting actions through the actor, receiving rewards, and updating the Q-networks and actor based on experiences stored in a replay buffer. These innovations make TD3 effective for precise continuous control in applications like autonomous guidance and cooperative UAV systems.
@@ III-A2 Artificial Potential Field Based Reward Structure
Let $a_{\text{opt}}$ represent the optimal action (direction) derived from the APF, and $a_{\text{agent}}$ represent the action taken by the drone in a given state. The reward function can be designed to encourage the agent to choose actions that align with the optimal direction. One way to measure this alignment is by using the cosine similarity between the two vectors:
$R=\frac{a_{\text{opt}}\cdot a_{\text{agent}}}{\|a_{\text{opt}}\|\|a_{\text{agent}}\|},$ (3)
[…]
Alternatively, a simpler formulation could penalise the agent based on the difference between the actions:
$R=-\|a_{\text{opt}}-a_{\text{agent}}\|,$ (4)
where the reward becomes smaller as the chosen action deviates further from the optimal direction.
@@ V-A1 Challenges in Height (Altitude) Estimation in Enclosed Environments
VI Methodology
@@ VI-A UAV and Sensor Suite
VI-B LIDAR-SLAM Altitude Fix
The downward-facing Red, Green, Blue, and Depth (RGBD) camera is utilised to correct for the predicted height drift. First, the median depth was calculated from five equidistance areas in the depth image. One of these was located in the centre of the image, and the other four were in the centre of each quadrant. The final depth was measured as the median of the five values and injected into the control loop.
@@ VI-B LIDAR-SLAM Altitude Fix
VI-C Twin Delayed deep Deterministic Policy Gradient
The TD3 scheme, as previously described, will be utilised to train the two DNNs responsible for providing guidance and task allocation. This scheme enhances the DDPG approach by incorporating three core innovations. The first is an additional critic network for Q-learning. The Q values generated by the two critic networks can be compared, allowing for the utilisation of the minimum Q value during the training process. This methodology effectively constrains the updates applied to the actor network, thereby mitigating the risk of overestimation. The scheme also uses delayed policy updates to stabilise critics before they can update the actor, preventing premature updates. The final core innovation is target policy smoothing, where noise is added to target actions to improve exploration and reduce overfitting.
Additional techniques were used to improve the training process for this implementation. The first is a prioritised experience replay, which ranks the experience’s quality, consisting of the state, next state, action, reward, and done. When the mini-batch is selected from the memory buffer, the random selection is weighted to be more likely to pick higher learning potential experiences. By focusing the learning using experiences that the agent found surprising or had significant errors, the agent will converge quicker and more effectively [14]. The TD error defines the learning potential, which is calculated using one of the critic networks. Secondly, to calculate the loss of the critic networks, the Huber loss was used rather than the usual mean squared error loss to protect the training from outliers.
The next sections will describe the architecture that will be used to make the decisions needed for the proposed tasks. In the TD3 scheme shown in Fig. 5, these networks play the part of the actor networks. The critic networks are the same as the actor networks but include a branch that inputs the actor’s output into the concatenation layer. Unlike the actor networks, the final output is just one Q-value. The target networks are copies of the main networks that stabilise the training process.
VI-D Guidance Deep Neural Network Architecture
The proposed DNN architecture for the actor network for the UAV perception component of our guidance TD3 DRL scheme integrates three input branches, each processing a different type of perception sensory data—depth images, LIDAR scans, and positional data—before combining them to predict the UAV’s motion commands to navigate complex environments.
VI-D1 Branch 1: Depth Image Processing
The first branch processes depth images captured from the UAV’s onboard camera. The depth image is passed through 8 convolutional layers [21], each using a LeakyReLU [26] activation function and batch normalisation [17] to stabilise the training process and improve convergence. The convolutional layers are divided into four blocks, each ending with a max-pooling operation to reduce the spatial dimensions and retain important features while providing translational invariance. The output of this branch is then flattened and passed through two fully connected (linear) layers, reducing the feature vector to a size of 128. This vector is further transformed using a tanh [6] activation function, preparing it for concatenation with the outputs from the other branches.
VI-D2 Branch 2: LIDAR Scan Processing
The second branch processes a 1D LIDAR scan, which provides a flat, radial representation of the surrounding obstacles. This input is passed through 5 one-dimensional convolutional layers, also using LeakyReLU activation and batch normalisation. The layers are organised into three blocks, each reducing the dimensionality of the input data while extracting spatial features relevant to obstacle proximity and distribution around the UAV. A final linear layer compresses the resulting feature map into a 128-dimensional vector, followed by a tanh activation, aligning the output with that of the first branch.
VI-D3 Branch 3: Positional Data Processing
The third branch takes a vector containing various positional parameters: the distance and bearing to the goal, the difference in altitude between the UAV and the goal, the relative height of the UAV, and the previous actions in the x, z, and yaw directions. This input is processed through a single fully connected layer with a tanh activation, resulting in a 128-dimensional vector. This approach allows the network to incorporate dynamic task-specific parameters directly, aiding the guidance decision-making process.
VI-D4 Fusion and Output Layer
The outputs from the three branches are concatenated into a single 384-dimensional vector. This combined representation is then passed through a series of 5 fully connected layers with LeakyReLU activations. These layers progressively refine the fused features, enabling the network to learn complex relationships between the depth, LIDAR, and positional information. A final linear layer with a tanh activation function outputs the predicted control commands, including the x-velocity, z-velocity, and yaw speed of the UAV. The tanh activation ensures that the output values remain bounded between -1 and 1, which is often desirable for controlling the UAV’s speed and orientation in a stable manner.
VI-D5 Discussion of Architectural Choices
The use of convolutional layers in the depth image and LIDAR branches allows the network to capture spatial hierarchies and patterns within the sensor data, which is crucial for understanding the environment’s structure. Batch normalisation aids in stabilising training by normalising the inputs of each layer, reducing internal covariate shifts. The LeakyReLU activation mitigates the risk of vanishing gradients in the early layers, ensuring more efficient learning. Tanh activation in the final layers helps ensure smooth control outputs, which is critical for the stability of UAV motion.
This multi-branch design enables the UAV to integrate depth-based obstacle detection, LIDAR-based spatial awareness, and task-specific positional data, resulting in a robust perception/guidance strategy suitable for GNSS-denied indoor environments.
The layout of the network can be seen in Fig. 4. The three separate branches that process the three inputs to the model are shown with the four 2D convolutional blocks, the three 1D convolutional blocks, and the fully connected layers for the data input. The layout of the FCN can be seen as combining and processing all the data together to produce the commands for the UAVs.
This network is trained using the TD3 algorithm shown in Fig. 5, where the network architecture shown in Fig. 4 is used as the actor and target actor networks. The critic networks are the same as the actor networks but add the action produced to the concatenation layer and produce a single Q-value.

### v2

@@ Abstract
This paper presents the first end‑to‑end framework that combines guidance, navigation, and centralised task allocation for multiple UAVs performing autonomous search‑and‑rescue (SAR) in GNSS‑denied indoor environments. A Twin Delayed Deep Deterministic Policy Gradient controller is trained with an Artificial Potential Field (APF) reward that blends attractive and repulsive potentials with continuous control, accelerating convergence and yielding smoother, safer trajectories than distance‑only baselines. Collaborative mission assignment is solved by a deep Graph Attention Network that, at each decision step, reasons over the drone–task graph to produce near‑optimal allocations with negligible on‑board compute. To arrest the notorious Z‑drift of indoor LiDAR‑SLAM, we fuse depth‑camera altimetry with IMU vertical velocity in a lightweight complementary filter, giving centimetre‑level altitude stability without external beacons. The resulting system was deployed on two 1 m‑class quad‑rotors and flight‑tested in a cluttered, multi‑level disaster mock‑up designed for the NATO‐Sapience Autonomous Cooperative Drone Competition. Compared with prior DRL guidance that remains largely in simulation, our framework demonstrates an ability to navigate complex indoor environments, securing first place in the 2024 event. These results demonstrate that APF‑shaped DRL and GAT‑driven cooperation can translate to reliable real‑world SAR operations.
@@ I Introduction
Key Contributions.
Unlike prior work that validates DRL controllers only in small-scale labs or simulations, we report the end-to-end real-world deployment of a cooperative DRL policy for autonomous SAR with multiple UAVs. Our main novelties are:
  1. 1.
    Extensive real‑world evaluation. We flew two medium-sized (1.0m at widest) quad‑rotors across a large, structurally complex disaster mock‑up. This paper details the lessons learned from tackling the sim-2-real problem in multi-drone guidance.
  2. 2.
    APF-referenced action-alignment reward for continuous velocity-level UAV control. We propose an APF-referenced shaping objective in which an APF-inspired scheme produces a step-wise reference action $a_{t}^{\star}$ in the same control space as the UAV commands, and the DRL agent is rewarded for aligning its executed action $a_{t}$ with this reference. The design includes obstacle-proximity mode switching (attractive vs. repulsive behaviour), reward clipping for training stability, and terminal overrides for goal completion and collision. We demonstrate that this reward enables stable training and real-world deployment of a continuous-velocity guidance policy in a cluttered indoor environment where a standard distance-to-goal reward did not produce a deployable controller under the same training constraints.
  3. 3.
    Centralised task allocation via Graph‑Attention DRL. We formulate the cooperative search as a one‑hop message‑passing problem and train GAT‑actor–critic that centrally assigns tasks while each UAV executes local low‑level policies.
  4. 4.
    Altitude‑drift suppression for indoor LiDAR‑SLAM. To mitigate the well‑known Z‑drift of 3‑D LiDAR SLAM indoors, we fuse a depth camera–derived altitude with the IMU’s vertical velocity via a lightweight complementary filter, recovering a drift‑free global $z$ estimate.
Taken together, these novelties enable, for the first time, robust, collaborative DRL‑guided SAR flight in complex, GPS‑denied indoor environments.
@@ II-A Deep Reinforcement Learning for Autonomous Guidance and Obstacle Avoidance
To address this inefficiency, researchers have incorporated expert knowledge into DRL training, enhancing performance through pre-training. Techniques such as Deep Q Learning from Demonstrations (DQfD) [13] and Deep Deterministic Policy Gradients from Demonstrations (DDPGfD) [56] leverage expert-labelled data to pre-train networks, reducing reliance on extensive reward engineering while preserving the benefits of DRL.
Artificial Potential Fields (APFs) are widely used in UAV navigation as structured priors for goal-seeking and obstacle avoidance. When combined with DRL, the literature largely falls into three recurrent integration patterns: (i) using APF to guide the policy during training, (ii) embedding APF quantities within the reward as a shaping signal, and (iii) treating APF as an auxiliary controller or planner that is fused with, or adapted by, a learned policy.
One pragmatic approach is to generate APF ”non-expert” actions that steer learning early on, whilst gradually reducing this influence to restore exploration. NPE-DRL [60] follows this philosophy in a discrete action space, whereas DDPG-APF [14] demonstrates an analogous idea under continuous control. In both cases, the APF acts as a curriculum-like scaffold: it provides a stabilising directional cue in the initial phases of training, and is annealed so that the final policy is not constrained by the hand-crafted field. Empirically, these approaches typically improve sample efficiency and reduce the frequency of catastrophic exploration during early training.
A second class of methods incorporates APF structure directly into the reward function, using potential-inspired terms to densify feedback beyond sparse goal completion rewards. Li and Wu [23] integrate APF-inspired obstacle avoidance with a line-of-sight tracking objective, combining task-driven terms with safety-oriented penalties and action costs. Dong et al. [5] similarly decompose reward into attractive, repulsive, and directional alignment components, thereby encouraging both progress and locally sensible avoidance. These formulations can be effective in guiding the policy towards safer behaviour, but they remain sensitive to the relative weighting of reward components and, in several instances, are paired with discretised manoeuvre sets (e.g., a finite set of yaw-rate and vertical velocity commands), which can limit fine control authority in confined environments.
A third integration pattern uses APF to bias or constrain actions at execution time. Shield-DDPG, for example, augments the action generation process with an APF-inspired attractive component and employs a safety “shield” that overrides unsafe actions [25]. Such designs can provide explicit constraint satisfaction, although the deployed behaviour is partly determined by the override logic and may differ from the learned policy. Related hybrid approaches treat APF as a planner whose behaviour is adapted by learning. PG-ITD3 combines a TD3 Variant with potential-field guidance and introduces mechanisms to mitigate classic APF failure modes (e.g., local minima and goal non-reachability), while also accounting for kinematic constraints [43]. In these frameworks, APF provides a strong inductive bias for navigation, while learning adjusts parameters or higher-level decisions to improve robustness.
This paper is most closely related to APF-guided and APF-shaped DRL methods, but differs in where the potential field is placed in the learning loop and in the targeted deployment setting. In contrast to approaches that treat APF as an external ”teacher” through action injection (e.g., NPE-DRL [60] and DDPG-APF [14]), we integrate APF directly into the guidance objective by shaping reward in the same continuous action space used for deployment. Relative to reward-only potential shaping, we emphasise that our formulation produces dense, physically meaningful feedback aligned with the commanded control variables, which is particularly important for indoor, GNSS-denied flight, where tight clearances and non-ideal dynamics demand smooth and responsive velocity and yaw-rate control. Finally, unlike safety-override schemes, our intention is for the learned policy to internalise the desired avoidance behaviour rather than routinely relying on an external action filter, improving consistency between training and fielded execution. We provide real-world experimental data to back up the
Talha et al. [54] used DRL to generate waypoints for a UAV in an AirSim SAR simulation, combining learned guidance with classical planners and detectors. In contrast, our controller outputs direct velocities and is validated in a cluttered indoor setting where fine control is essential, rather than the comparatively forgiving outdoor scene they considered.
A persistent weakness in the literature is the scarcity of real-world indoor tests. Phueakthong et al. [41] opted for a position-based (2‑D) controller to limit aggressive manoeuvres, feeding a 26‑zone LiDAR scan and target pose into an FCN. Their platform (1.9 kg, 0.1 m/s) successfully rounded a corner, but the action/state spaces were deliberately constrained. We expand both: yaw and $z$ velocity are controlled, a depth image and a denser LiDAR scan are processed, and our 4 kg UAVs fly at up to 1.0 m/s in a far more complex layout.
Other efforts have begun to probe sim‑to‑real transfer. Using PPO, [18] trained a micro‑drone to orbit a column and injected Gaussian noise to harden the policy against sensor errors, improving real-world transfer in a small arena. While valuable, that study still operates at a miniature scale and with a limited task.
Real-world DRL guidance is more common in outdoor settings. He et al. [12] emphasise explainability and demonstrate tree‑traversal with a 2‑D flight envelope. Their contribution complements ours—interpretability versus operational breadth—but does not align with the scope of our indoor, multi-UAV deployment.
In summary, prior work either remains in simulation, restricts dynamics/action spaces, or tests at a small scale. Our study addresses this gap with full‑scale, multi‑agent DRL guidance in a GNSS-denied indoor environment, extensive real-world trials, and centralised GAT-based task allocation.
@@ II-B Deep Reinforcement Learning for Graph-Based Task Allocation
Alshaboti and Baroudi[1] compare two distributed multi-robot task-allocation (MRTA) strategies—a fuzzy-inference auction and an adaptive multi-threshold rule—in dynamic, single-task/single-robot settings where tasks arrive online and multiple objectives (distance, load balance, “quality” match) must be traded off. Their Mamdani FIS compresses distance, load, and quality-gap metrics into a scalar bid, while the threshold scheme allows each robot to decide locally. Simulations with Khepera III robots in Webots show that the threshold method cuts the travelled distance by roughly 10% but suffers from slightly worse load balance, whereas the auction achieves tighter balancing with similar quality of satisfaction. Both approaches, however, presuppose reliable communication, relatively simple robots, and GPS/landmark localisation, and they do not model rich perception or high-dimensional action spaces. In contrast, our work allocates tasks centrally via a graph-attention actor–critic that reasons over dense edge costs and heterogeneous node states, and—critically—validates the policy on full-scale UAVs in a GNSS-denied indoor environment rather than in simulation. This highlights a complementary trajectory: moving from scalar bids and hand-crafted fuzzy rules toward learned, end-to-end differentiable allocators embedded in a real SAR pipeline.
@@
Ryu et al.[44] propose HAMA, a multi‑agent actor–critic architecture that stacks hierarchical graph attention layers—separate inter‑agent and inter‑group attentions—to encode relationships and enable scalable, transferable policies in mixed cooperative–competitive games. They train under CTDE and show superior performance and transfer across predator–prey benchmarks (3v1, 3v3, “more‑the‑stronger”) compared to MADDPG, MAAC, and heuristics, while also interpreting strategies via learned attention weights. In contrast, our allocator is centralised (single forward pass assigns tasks to all UAVs) and is validated on full-scale robots in a GNSS-denied indoor SAR scenario, addressing sim-to-real, rich perception, and safety constraints that are absent from purely simulated domains.
@@ III-A1 Twin Delayed Deep Deterministic Policy Gradient
TD3[9] refines DDPG[26] for continuous control by (i) clipped double Q‑learning, which updates from the smaller of two critic estimates to curb overestimation, (ii) delayed actor updates that let critic values stabilise, and (iii) target‑policy smoothing, adding small noise to target actions to regularise Q targets. The agent stores transition tuples in a replay buffer, updates both critics and the actor using mini-batches in an off-policy loop, and outputs smooth, high-precision control signals, making TD3 well-suited to the cooperative UAV guidance task studied here.
@@ V-A1 Challenges in Height (Altitude) Estimation in Enclosed Environments
VI Problem Formulation
We consider a cooperative UAV system composed of two learned decision-making modules: (i) a continuous-control guidance policy executed onboard each UAV for collision-aware navigation; and (ii) a centralised task allocation policy executed at the ground station for assigning tasks to multiple UAVs. The guidance policy handles short-horizon motion selection under partial observability, whilst the allocator operates at a lower frequency to coordinate high-level tasking.
VI-A Guidance as a Partially Observable Markov Decision Process
Each UAV navigation episode is modelled as a partially observable Markov decision process (POMDP). At time-step $t$, the UAV receives a multi-modal observation
$o_{t}=\big(I_{t}^{\text{depth}},\,\ell_{t}^{\text{LiDAR}},\,p_{t}\big),$ (5)
where $I_{t}^{\text{depth}}$ is the forward-facing depth image, $\ell_{t}^{\text{LiDAR}}$ is a 1-D LiDAR sweep, and $p_{t}$ is a compact positional state containing goal-related quantities (e.g., range/bearing and altitude error) together with recent action history. The policy outputs a continuous bounded action
$a_{t}=[v_{x,t},\,\omega_{t},\,v_{z,t}]\in[-1,1]^{3},$ (6)
corresponding to forward velocity, yaw-rate, and vertical velocity commands. The UAV then transitions to a new state via the environment dynamics and receives an immediate reward $r_{t}$.
APF-derived optimal action.
We compute an APF-inspired “optimal” reference action $a_{t}^{\star}$ from goal geometry and obstacle proximity. When the UAV is beyond a $1\,\mathrm{m}$ proximity threshold from obstacles, the reference is governed by attractive terms:
$v_{x,t}^{\star}$ $=\mathrm{clip}(\cos(\theta_{t}),\,0.0,\,1.0),$ (7)
$\omega_{t}^{\star}$ $=\mathrm{clip}\!\left(\theta_{t}\cdot\frac{3}{\pi},\,-1.0,\,1.0\right),$ (8)
$v_{z,t}^{\star}$ $=\mathrm{clip}\!\left(2\,(z_{\text{goal}}-z_{t}),\,-1.0,\,1.0\right),$ (9)
where $\theta_{t}$ is the heading error between the UAV and the target direction.
When an obstacle is within $1\,\mathrm{m}$, we switch to a repulsive mode. The forward reference velocity is set to move away from obstacles according to the obstacle bearing $\theta_{\text{obs}}$:
$v_{x,t}^{\star}=-0.5\cos(\theta_{\text{obs}}),$ (10)
and the yaw-rate reference follows a piecewise turning function $\omega^{\star}(\theta_{\text{obs}})$ that turns away from the obstacle depending on its relative angle.
Reward.
The guidance reward is designed to encourage the UAV to align its executed action with the APF reference action. The base reward is
$r^{\text{base}}_{t}=1-\sum_{i\in\{x,\omega,z\}}\big(a_{t,i}-a^{\star}_{t,i}\big)^{2},$ (11)
which is then clipped for stability:
$r_{t}=\mathrm{clip}(r^{\text{base}}_{t},\,-1.5,\,1.0).$ (12)
In addition, terminal outcomes override the clipped shaping reward: reaching a goal yields $r_{t}=R_{\text{goal}}=2$, whilst colliding with an obstacle yields $r_{t}=R_{\text{collision}}=-2$.
Objective and termination.
The guidance policy $\pi_{\phi}(a_{t}|o_{t})$ is trained to maximise the expected discounted return
$J(\pi_{\phi})=\mathbb{E}\left[\sum_{t=0}^{T-1}\gamma^{t}r_{t}\right],$ (13)
where episodes terminate on goal completion, collision, or a fixed horizon $T$.
VI-B Task Allocation as a Centralised Graph Decision Process
Task allocation is cast as a centralised RL problem operating at decision epochs $t$ over a fully-connected graph $G_{t}=(V_{t},E_{t})$ comprising UAV nodes and task nodes. Node features encode completion/occupancy indicators, and edge attributes encode collision-free shortest-path distances between entities. The allocator outputs an assignment matrix
$a_{t}=[a_{nk}]\in\{0,1\}^{N\times(M+1)},$ (14)
where $a_{nk}=1$ assigns UAV $n$ to task $k$ (or to an “idle” task $k=0$), subject to $\sum_{k}a_{nk}=1$ for each UAV.
The allocation reward credits productive completion whilst penalising inefficiency and collisions:
$r_{t}=\Delta C_{t}-\lambda_{d}\Delta D_{t}-\lambda_{c}C_{t},$ (15)
where $\Delta C_{t}$ is the number of newly completed tasks, $\Delta D_{t}$ is the incremental distance flown above a baseline, and $C_{t}$ indicates collision events. Episodes terminate when all tasks are complete or a fixed horizon is reached.
VII Methodology
@@ VII-A UAV and Sensor Suite
VII-B LIDAR-SLAM Altitude Fix
To fix this problem, the most obvious solution would be to use a laser range finder or a radar altimeter. We were unable to use these as a solution because the parts were not on the approved competition list, so another solution was needed. In the competition equipment list, a second Intel RealSense was allowed, so we moved on to using this as our range finder.
To utilise the RGB-D camera, we need to find a robust way to measure the correct distance. First, the median depth was calculated from five equidistant areas in the depth image. One of these was located in the centre of the image, and the other four were in the centre of each quadrant. The final depth was measured as the median of the five values and injected into the control loop.
This method was used to improve the accuracy of the measured distance. The depth camera is a noisy sensor, with values that can fluctuate widely. So by using five areas and using the median values, we factor in outliers when taking a measurement. The camera’s wider field of view also provides a clearer view of what is beneath the drone as it flies, compared to a more typical laser rangefinder. This means that when negotiating a ledge, we can be more certain when the edge is cleared.
@@ VII-B LIDAR-SLAM Altitude Fix
VII-C Guidance Deep Neural Network Architecture
The actor network in our TD3 guidance scheme fuses three perception streams—depth images, LiDAR scans and positional/task variables—to output velocity commands for the UAV.
VII-C1 Branch 1: Depth Images
Depth frames are processed by eight 2‑D convolutional layers grouped into four blocks, each followed by max‑pooling. LeakyReLU activations and batch normalisation stabilise training [22, 29, 17]. The feature map is flattened and passed through two linear layers to yield a 128‑D vector, finished with a tanh activation.
VII-C2 Branch 2: LiDAR Scan
A 1‑D LiDAR sweep is handled by five 1‑D convolutional layers arranged in three blocks, again with LeakyReLU and batch normalisation. A final linear layer compresses the representation to 128 dimensions and applies tanh.
VII-C3 Branch 3: Positional State
A compact vector (goal range and bearing, altitude error, relative height, and previous $x$, $z$ and yaw actions) is mapped to a 128‑D embedding through a single fully connected layer with tanh activation.
VII-C4 Fusion and Output
The three 128‑D embeddings are concatenated (384‑D) and refined by five fully connected layers with LeakyReLU. A final linear layer with tanh outputs bounded xx‑velocity, zz‑velocity and yaw‑rate commands (-1 to 1), aiding stable control.
VII-C5 Training Context
This network serves as the actor (and target actor) in TD3 (Fig. 5); the critics share the same architecture but concatenate the action at the fusion stage and regress a single Q‑value.
VII-C6 Design Rationale
Convolutions capture spatial structure in depth and LiDAR inputs, while batch normalisation and LeakyReLU mitigate internal covariate shift and vanishing gradients. The final tanh ensures smooth, bounded control signals, which is advantageous for precise indoor flight. The multi-branch design enables the exploitation of complementary sensing modalities without requiring hand-crafted feature engineering (Fig. 4).